OpenStudy (anonymous):

0.05X+0.25 Y=11 0.15+0.05 Y = 12 SOLVE BY ELIMINATION METHOD

OpenStudy (anonymous):

do you mean 0.15X + 0.05Y = 12? If so ,multiply that whole equation by 5 so you get 0.75X + 0.25 = 60 now subtract that first equation to give 0.7X=48

OpenStudy (anonymous):

sorry i made a mistake in my 2nd equation (4th line), its meant to be 0.75X + 0.25Y = 60 (i forgot the y)

OpenStudy (anonymous):

If you mean 0.15x + 0.05y = 12, we'll have: 0.15x + 0.75y = 33 0.15x + 0.05y = 12 0.70y = 21 y = 21/0.70 or 2100/70 or 30 0.15x + 0.05y = 12 0.15x + 0.05(30) = 12 0.15x + 1.5 = 12 0.15x = 10.5 x = 105/15 or 7

OpenStudy (anonymous):

SO WHATS THE ANSWER IM KIND OF CONFUSED

OpenStudy (amistre64):

0.05x+0.25 y =11 ; multiply by -1 0.15x+0.05 y = 12 -0.05x-0.25 y =-11 0.15x+0.05 y = 12 ------------------- .10 x = 1 x = 10 0.15(10)+0.05 y = 12 1.5+0.05 y = 12 0.05 y = 12-1.5 y = 10.5/.05 = 1050/5 = 210

OpenStudy (amistre64):

0.05(10)+0.25(210) =11 .5 + 52.5 ... got it wrong .... bummer

OpenStudy (anonymous):

x = 70 and y = 30

OpenStudy (amistre64):

70 and 30 are right :)

OpenStudy (anonymous):

0.05X+0.25Y = 11 0.15X+0.05Y = 12 5x + 25y = 1100 15x + 5y = 1200

OpenStudy (anonymous):

Oh. I got the 10.5/0.15 wrong. It must be 1050/15 or 70, as La-MrC said.

OpenStudy (amistre64):

i saw the numbers wrong .... prolly cause of the decimals :)

OpenStudy (anonymous):

Yeah. Decimals can really be confusing in the eyes sometimes.